Find the orthogonal trajectories of the family of parabolas x = cy²
1
Expert's answer
2019-09-04T09:13:09-0400

Let "(x,y=Y)" be the intersection point of the curves

"y=\sqrt{x/c} \quad \bot \quad Y=Y(x), \quad x>0, \quad y>0."

Then

"Y'=-1/y' \quad \lrArr \quad Y'=-2\sqrt {cx}."

But "c=\frac{x}{y^2}." Therefore

"Y'=-\frac{2x}{Y} \quad \text{or} \quad YdY=-2xdx \quad \text{or} \quad \int YdY=-\int 2xdx."

Then

"Y^2+2x^2=C"

We get the family of ellipses finally:

"\frac{x^2}{a^2} + \frac{y^2}{2a^2}=1".